OpenStudy (anonymous):

You have a coupon good for $5 off the price, p, of any large pizza. You also get a 15% discount on any pizza, if you show your student ID. How much more would you pay for a large pizza, if the cashier applies the coupon first (rather than the 15% discount)?

OpenStudy (ineedhelplz):

Ok. well you need two simple equations. FOR USING THE COUPON FIRST: (p-5)(0.85) FOR USING THE DISCOUNT FIRST: (0.85)p)-7 so now, plug in any price for the pizza. for example, let's say the pizza was originally $20. (20-5)=15 times .85=$12.75<<<

OpenStudy (anonymous):

Where did you get .85 from? The rest makes sense to me

OpenStudy (anonymous):

@Ineedhelplz

OpenStudy (anonymous):

And where did you get -7 in the second equation?

OpenStudy (ineedhelplz):

x is the price of the pizza. if coupon is applied before: (x-5)0.85=0.85x-0.85*5 if it is applied after: 0.85x-5 Therefore, the difference between the two is 0.15*5=0.75

OpenStudy (ineedhelplz):

Is that better?

OpenStudy (anonymous):

Yes. So the answer is 75 cents?

OpenStudy (ineedhelplz):

Yes

OpenStudy (ineedhelplz):

Good Job :)

OpenStudy (anonymous):

Oh okay. Thank you!

OpenStudy (ineedhelplz):

You're welcome